OpenStudy (anonymous):

(a) 4 dm expressed in mm mm (b) 8 h 10 min expressed in s s (c) 19 g expressed in µg µg (d) 0.55 km expressed in cm cm (e) 0.700 mg expressed in g g (f) 452 µm expressed in cm cm (g) 30 km/h expressed in m/s m/s

OpenStudy (praxer):

sorry #typo the Left side unit to 1

OpenStudy (anonymous):

thank you, but idk. i always have trouble with finding out how to convert all of this. I know we have to use KHDUDCM, but i am always stuck when it comes to using this in a problem.

OpenStudy (praxer):

Okay! for length chart 1. https://www.google.co.in/webhp?sourceid=chrome-instant&ion=1&espv=2&ie=UTF-8#q=30+km%2Fh+to+m%2Fs 2. for time always remember that 1 min =60 seconds and 1 hour= 60 mins which means 1 hour = 60*60=3600 seconds. 3. For weight remember that, 1 kg = 1000 gram and 1 micro gram =1*10^-6 gram 4. 1 km = 1000 cm 5.1000mg = 1g 6.1 micrometer = 1*10^-6meter 7.1km/hr= 5/18(m/sec) ref; http://www.mathsisfun.com/length-conversion.html

OpenStudy (praxer):

So as for unit conversion problem you should start doing more and more of the conversions so that you can remember the charts. Starting with the basic will help the most. Like 1m =1000mm 1000m=1km 1g=1000mg 1kg=1000g http://www.math4ged.com/wp-content/uploads/2011/07/ged-math-metric-table2.png you may find this helpful kilometer hectometer dekameter meter(U) decimeter centimeter milimeter 1 | 10 | 100 | 1000 |10000 | 100000 |1000000 it is your KHDUDCM

OpenStudy (anonymous):

thank you so much!

OpenStudy (praxer):

$$\huge \color{red}{ Welcome \\\\\ddot\smile} $$